# Detecting AI-Generated Text: Factors Influencing Detectability with Current Methods

### v1

@@ 3.2.2 Stylistic Analysis
Also working in the domain of news articles, ? (?) compare and contrast the linguistic patterns observed in human-generated text and text generated by the Llama family of LLMs. Using articles from the New York Times as human text, they prompt four different Llama models to generate news articles based on the headline and the first three words of the lead paragraph. Linguistic analysis reveals that the AI generated text has a more restricted vocabulary, uses fewer adjectives and more symbols and numbers, includes fewer words associated with negative emotions, and has a higher propensity for male pronouns. However, the authors do not evaluate the usefulness of these features in a detector, and it is not clear how well these features would generalize across different model families (e.g. GPT) and domains.
@@ 3.3 Language Model-Based Classification
? (?) use a RoBERTaBASE model as their pretrained language model, and then fine-tune a multilayer perceptron classifier on top, keeping the RoBERTa weights frozen. They also experiment with a T5 model (?), which they train as a sequence-to-sequence model that outputs either ‘positive’ or ‘negative’, given the text as input. They test these two classification models on datasets of text generated by GPT-2 and GPT-3.5 and find that the T5-based model is generally superior; however, both models outperform the OpenAI GPT-detector and zero-shot learning with GPT-2.
@@ 4 AIGT Datasets
OpenGPTText (?) English Websites 30K
@@
As described in the previous section, there are three main methods for the detection of AI-generated text (AIGT): watermarking, statistical and stylistic analysis, and the use of pre-trained language models (LMs). The detection of a watermark requires knowledge of the watermark extraction algorithm; beyond that, no extra data are needed. However, the other two broad detection methods require data in order to learn the patterns that distinguish AIGT from human-written text – ideally, a dataset where the positive examples were generated by the AI model we wish to detect, and in a setting as close as possible to how we expect to encounter these texts in the real world. Previous work has indicated that the most effective detectors are trained on data from the same domain (news articles, social media posts, academic essays, etc.), language (English, Chinese, French, etc.) and model settings (decoding algorithm, prompt, length of output, etc.) as the test data. At the same time, the research also shows that for maximal generalizability and robustness, it is essential to train a detector on a wide variety of data so that it is not overfit to one very narrow range of data samples. Therefore, for any given application, it is important to select appropriate data to first train the detector, and also to test the detector for its accuracy: “The effectiveness of black-box detection models is heavily dependent on the quality and diversity of the acquired data.” (?)
@@
As mentioned above, another important factor is the domain of the dataset. We observe the presence of expected domains such as news, academic writing, and essays. Social media, while highly interesting from a practical perspective, is not well-represented, potentially due to the difficulty of detecting AIGT in short texts, or of simulating realistic social media posts for dataset creation. Some datasets that focus more on casual, online settings with multiple authors include: HC3 (?), M4 (?), In-the-wild (?), MixSet (?), OpenGPTText (?), HC-Var (?), Real/Fake Tweets by ? (?), and COVID-19 Related AI-Generated Tweets by ? (?). A fairly large proportion of the datasets include multiple domains. For example, the M4 dataset (?) is a large multi-purpose dataset containing seven different languages, and multiple domains and generators, including state-of-the-art LLMS like GPT-4. Not surprisingly, many datasets with AI-generated content were created for the task of misinformation detection. We further note that within this broad category, there are multiple sub-categories (e.g., fake news, social media misinformation, etc.).
@@
Instead, most datasets are generated by first defining a “parent” or “anchor” dataset of human-generated text, and then artificially generating parallel AIGT text. For example, in the news domain, researchers might start with a corpus of news stories. They then feed the headline (or the headline plus the first sentence) into an LLM and ask the LLM to write the rest of the article. In this fashion, they can generate a parallel dataset of human- and AI-written articles for the same set of headlines. Similarly, in the domain of question-answering, if the researcher has a dataset of questions with human-written answers, they can ask the AI to answer the same questions. Some generation methods include more specific style prompting to mimic the human dataset, as in “Write a news article in the style of the New York Times.” A more challenging detection scenario is controlled generation, where the LLM is more constrained by the parameters of an existing human-written text, as in the case of summarization or paraphrasing. Datasets that tackle this issue include CHEAT (?), GPABench2 (?), tum-nlp/IDMGSP (?), HC3-SI (?), and OpenGPTText (?). An even more challenging scenario for AIGT detection is mixcase, or documents that contain a mix of human and AI writing. Some of the datasets that include this type of data are: MixSet (?), GPAbench2 (?), and SeqXGPT-Bench (?). Such mixcase scenarios are especially critical in the domain of misinformation detection where a documents containing a mixture of true and false statements pose a significant challenge for detection algorithms. Such scenarios include: perturbing just a few facts in real human-written articles, as in F3 (?) and ODQA (?), removing and adding negations, as in AI-Generated News by ? (?), finding and replacing the most salient sentence in the story, as in PropaNEWS (?), or mixing a real and a false story in one article, as in Human-Written and AI-Generated Fake News by ? (?). Further, generating and editing a story in a multiple-step dialog with the model, as done in e.g., AI-Generated Fake News by ? (?) and Human-Written and AI-Generated Fake News by ? (?), may produce even more sophisticated fake news, but would require more human resources. All these methods of generating text allow for precise, controlled research experiments on detecting AIGT; however, it is unclear how well they represent the artificially generated text that actually exists on the internet.
@@ 5.1 Properties of the Generating LLM
Generally speaking, increasing the size (i.e. number of trained parameters) of a language model is one way to improve the quality of the generated text. Unsurprisingly then, larger language models (e.g. GPT-4) are more difficult to detect. ? (?) introduce the AI Detectability Index (ADI), an experimentally motivated index to quantify the detectability of LLMs. Based on their observation that newer, larger language models have statistical signatures (e.g. perplexity and negative-log curvature signals) approaching that of human distributions, they assign higher ADI to larger models. ? (?) show that this property holds true even when the same generating model is fine-tuned for detection (i.e. the detector has the same capacity for representing language as the generator). Experimentally, they observe that detectability is related to model size by a power law, meaning that detection accuracy decreases linearly as the number of model parameters increases exponentially.
Secondly, the decoding strategy of the generating model affects detectability. Recall from Section 2.1 that a decoding strategy refers to the method of choosing the next word to generate, given the probability distribution over possible choices. Top-$k$ sampling selects among a fixed number of choices, whereas nucleus sampling allows more possible choices when the model is uncertain. ? (?) observe that detection accuracy decreases significantly when training data is generated using a different decoding strategy than test data (e.g. a decrease of 21% is observed when the detector is trained on top-$k$ decoding output and tested on nucleus sampling output compared to in-distribution results). In general, nucleus sampling output is the most difficult to detect, and detectors trained on nucleus-sampled data also generalize the best across other sampling methods. ? (?) show that even within the same decoding strategy, changing the decoding parameters between training and inference time can significantly impact detectability. For instance, when using nucleus sampling, the recall of a statistical detector on AIGT drops by 13% by changing the probability threshold from 0.96 to 0.8. Likewise, for top-$k$ sampling, recall falls by 56.4% by changing the $k$ from 40 to 160. ? (?) also highlight the potential difficulty of detecting text output by a generative discriminator decoding strategy. In this setup, a smaller auxiliary language model is used to guide the larger language model toward generating text with a specific attribute (e.g. positive sentiment). Most detection methods do not seem to consider the possibility of generative discriminator decoders.
@@ 5.6 Adversarial Attacks
5.7 Off-the-Shelf Evasion Tools
The literature on adversarial attacks helps to inform off-the-shelf evasion tools. In real-life applications of LLMs, many users wish to hide the fact that their writing is AI-generated. For example, students who submit AI-generated content can face academic misconduct charges, and AI-generated content on social media can be automatically blocked or censored. In response to AIGT detection tools, many off-the-shelf evasion tools are now available as well, and some of these are summarized in Table 4.
Name URL Language Evasion technique(s)
QuillBot quillbot.com 23 lang. Paraphrase (many styles); Synonym replacement
HideMyAI hidemy.ai 4 lang. Paraphrase (fluency match); Watermark removal
Undetectable AI undetectable.ai 4 lang. Paraphrase
Content at Scale contentatscale.ai English Paraphrase
Netus AI netus.ai English Paraphrase (voice-tuned); Watermark removal
GPTInf gptinf.com English Paraphrase
Table 4: Some of the available off-the-shelf evasion tools.
All the tools use paraphrasing attacks in some way to rewrite the text completely while maintaining the original message. Some tools allow more user control over the output, such as paraphrasing in a specific style (e.g., creative versus academic), or controlling the output fluency level. One tool claims to paraphrase while maintaining the original voice of the human author based on past writing samples. In other cases, the user has no control over the style, voice, or fluency of the output. Because these are commercial products, there is not much visibility into the underlying method, but we can assume they are using an unknown in-house language model as the paraphraser. The paraphrasing model is designed to be undetectable by the existing off-the-shelf detectors, likely by increasing variation in sentence structure and vocabulary choice, but new training data could be obtained from the evasion tools to fine-tune existing detection methods. Since paraphrasers are just unseen language models, the literature on detection methods is still applicable to this type of AIGT.
Some of the evasion tools also claim to remove watermarks, but do not provide any details about how this is verified. Since synonym replacement and paraphrasing are known to remove the primary type of watermark (i.e. one based on a red-green vocabulary list), evasion tool providers might be assuming their methods will remove any watermark, if it exists, without being able to detect the watermark themselves.

### v2

@@ 3.2.2 Stylistic Analysis
Also working in the domain of news articles, ? (?) compare and contrast the linguistic patterns observed in human-generated text and text generated by the Llama family of LLMs. Using articles from the New York Times as human text, they prompt four different Llama models to generate news articles based on the headline and the first three words of the lead paragraph. Linguistic analysis reveals that the AI generated text has a more restricted vocabulary, uses fewer adjectives and more symbols and numbers, includes fewer words associated with negative emotions, and has a higher propensity for male pronouns. However, the authors do not evaluate the usefulness of these features in a detector, and it is not clear how well these features would generalize across different model families (e.g., GPT) and domains.
@@ 3.3 Language Model-Based Classification
? (?) use a RoBERTaBASE model as their pretrained language model, and then fine-tune a multilayer perceptron classifier on top, keeping the RoBERTa weights frozen. They also experiment with a T5 model (?), which they train as a sequence-to-sequence model that outputs either ‘positive’ or ‘negative’, given the text as input (?). They test these two classification models on datasets of text generated by GPT-2 and GPT-3.5 and find that the T5-based model is generally superior; however, both models outperform the OpenAI GPT-detector and zero-shot learning with GPT-2.
@@ 4 AIGT Datasets
OpenLLMText (?) English Websites 30K
@@
As described in the previous section, there are three main methods for the detection of AI-generated text (AIGT): watermarking, statistical and stylistic analysis, and the use of pre-trained language models (LMs). The detection of a watermark requires knowledge of the watermark extraction algorithm; beyond that, no extra data are needed. However, the other two broad detection methods require data in order to learn the patterns that distinguish AIGT from human-written text – ideally, a dataset where the positive examples were generated by the AI model we wish to detect, and in a setting as close as possible to how we expect to encounter these texts in the real world. Previous work has indicated that the most effective detectors are trained on data from the same domain (news articles, social media posts, academic essays, etc.), language (English, Chinese, French, etc.) and model settings (decoding algorithm, prompt, length of output, etc.) as the test data. At the same time, the research also shows that for maximal generalizability and robustness, it is essential to train a detector on a wide variety of data so that it is not overfit to one very narrow range of data samples. Therefore, for any given application, it is important to select appropriate data to first train the detector, and also to test the detector for its accuracy: “The effectiveness of black-box detection models is heavily dependent on the quality and diversity of the acquired data” (?).
@@
As mentioned above, another important factor is the domain of the dataset. We observe the presence of expected domains such as news, academic writing, and essays. Social media, while highly interesting from a practical perspective, is not well-represented, potentially due to the difficulty of detecting AIGT in short texts, or of simulating realistic social media posts for dataset creation. Some datasets that focus more on casual, online settings with multiple authors include: HC3 (?), M4 (?), In-the-wild (?), AuTexTification (?), MixSet (?), OpenLLMText (?), HC-Var (?), Real/Fake Tweets by ? (?), and COVID-19 Related AI-Generated Tweets by ? (?). A fairly large proportion of the datasets include multiple domains. For example, the M4 dataset (?) is a large multi-purpose dataset containing seven different languages, and multiple domains and generators, including state-of-the-art LLMS like GPT-4. Similarly, the RAID benchmark (?) contains generations from eight different domains, 11 models, and four different decoding strategies. Not surprisingly, many datasets with AI-generated content were created for the task of misinformation detection. We further note that within this broad category, there are multiple sub-categories (e.g., fake news, social media misinformation, etc.).
@@
Instead, most datasets are generated by first defining a “parent” or “anchor” dataset of human-generated text, and then artificially generating parallel AIGT text. For example, in the news domain, researchers might start with a corpus of news stories. They then feed the headline (or the headline plus the first sentence) into an LLM and ask the LLM to write the rest of the article. In this fashion, they can generate a parallel dataset of human- and AI-written articles for the same set of headlines. Similarly, in the domain of question-answering, if the researcher has a dataset of questions with human-written answers, they can ask the AI to answer the same questions. Some generation methods include more specific style prompting to mimic the human dataset, as in “Write a news article in the style of the New York Times.” A more challenging detection scenario is controlled generation, where the LLM is more constrained by the parameters of an existing human-written text, as in the case of summarization or paraphrasing. Datasets that tackle this issue include CHEAT (?), GPABench2 (?), tum-nlp/IDMGSP (?), HC3-SI (?), and OpenLLMText (?). An even more challenging scenario for AIGT detection is mixcase, or documents that contain a mix of human and AI writing. Some of the datasets that include this type of data are: MixSet (?), GPAbench2 (?), and SeqXGPT-Bench (?). Such mixcase scenarios are especially critical in the domain of misinformation detection where a documents containing a mixture of true and false statements pose a significant challenge for detection algorithms. Such scenarios include: perturbing just a few facts in real human-written articles, as in F3 (?) and ODQA (?), removing and adding negations, as in AI-Generated News by ? (?), finding and replacing the most salient sentence in the story, as in PropaNEWS (?), or mixing a real and a false story in one article, as in Human-Written and AI-Generated Fake News by ? (?). Further, generating and editing a story in a multiple-step dialog with the model, as done in e.g., AI-Generated Fake News by ? (?) and Human-Written and AI-Generated Fake News by ? (?), may produce even more sophisticated fake news, but would require more human resources. All these methods of generating text allow for precise, controlled research experiments on detecting AIGT; however, it is unclear how well they represent the artificially generated text that actually exists on the internet.
@@ 5.1 Properties of the Generating LLM
Generally speaking, increasing the size (i.e., number of trained parameters) of a language model is one way to improve the quality of the generated text. Unsurprisingly then, larger language models (e.g., GPT-4) are more difficult to detect. ? (?) introduce the AI Detectability Index (ADI), an experimentally motivated index to quantify the detectability of LLMs. Based on their observation that newer, larger language models have statistical signatures (e.g., perplexity and negative-log curvature signals) approaching that of human distributions, they assign higher ADI to larger models. ? (?) show that this property holds true even when the same generating model is fine-tuned for detection (i.e., the detector has the same capacity for representing language as the generator). Experimentally, they observe that detectability is related to model size by a power law, meaning that detection accuracy decreases linearly as the number of model parameters increases exponentially.
Secondly, the decoding strategy of the generating model affects detectability. Recall from Section 2.1 that a decoding strategy refers to the method of choosing the next word to generate, given the probability distribution over possible choices. Top-$k$ sampling selects among a fixed number of choices, whereas nucleus sampling allows more possible choices when the model is uncertain. ? (?) observe that detection accuracy decreases significantly when training data is generated using a different decoding strategy than test data (e.g., a decrease of 21% is observed when the detector is trained on top-$k$ decoding output and tested on nucleus sampling output compared to in-distribution results). In general, nucleus sampling output is the most difficult to detect, and detectors trained on nucleus-sampled data also generalize the best across other sampling methods. ? (?) show that even within the same decoding strategy, changing the decoding parameters between training and inference time can significantly impact detectability. For instance, when using nucleus sampling, the recall of a statistical detector on AIGT drops by 13% by changing the probability threshold from 0.96 to 0.8. Likewise, for top-$k$ sampling, recall falls by 56.4% by changing the $k$ from 40 to 160. ? (?) also highlight the potential difficulty of detecting text output by a generative discriminator decoding strategy. In this setup, a smaller auxiliary language model is used to guide the larger language model toward generating text with a specific attribute (e.g., positive sentiment). Most detection methods do not seem to consider the possibility of generative discriminator decoders.
@@ 7 Challenges and Future Directions
Acknowledgments
This project was funded by the Canadian Department of National Defence, Annex Number: ER-VAL 23-085. The authors thank Zachary Devereaux for insightful feedback and discussion.